two forces of 10N and 6N act upon a body . the direction of the force is unknown. the resultant force on the body may be
1
Expert's answer
2018-07-19T15:44:08-0400
F_max=10+6=16 N.
F_min=10-6=4 N.
There is only one possible answer is 15 N.